The sum of two numbers is 89. Adding four times the first number to the second number gives 155. Whats the right way to express this equation?
Answer by checkley71(8403)

X+Y=89 OR Y=89-X
4X+Y=155
4X+(89-X)=155
4X+89-X=155
3X=155-89
3X=66
X=66/3
X=22 ANSWER FOR THE SMALLER NUMBER.
22+Y=89
Y=89-22
Y=67 ANSWER FOR THE LARGER NUMBER.
PROOF:
4*22+67=155
88+67=155
155=155